A 15-kilowatt motor is used to hoist an 800 kg bucket of concrete to the twentieth floor of a building under construction to height of 90 m. assuming no energy is lost along the process, how much time is required?


1
Expert's answer
2021-01-22T11:41:56-0500

Upward force needed =bucket's weight =(mg)


800"\u00d7" 9.8=7,840N


Power available (P)=1.0 × 15×103W=1.5×104W


Since P="\frac{W} {t} =\frac{Fs} {t}"


Therefore ;"t=\frac{7,840×90} {1.5×10^4}"


=470.4 seconds